OpenStudy (anonymous):

Is (8, -1) a solution of the system, 2x + 3y = 13 and x - y = 9?

OpenStudy (aroub):

(8,-1) 2x+3y=13 2(8)+3(-1)=13 16-3=13 13=13 x-y=9 8+1=9 So, yes it is a solution.

OpenStudy (anonymous):

Ok Thanks Both of you guys

OpenStudy (aroub):

You're welcome =)